OpenStudy (anonymous):

x^4-5x^2+4=0 find the real or imaginary solutions by factoring

OpenStudy (anonymous):

Did you find the imaginary roots of the last one? Or just need help w/ this one?

OpenStudy (anonymous):

need help with this one

OpenStudy (anonymous):

Factoring x4-5x2+4 The first term is, x4 its coefficient is 1 The middle term is, -5x2 its coefficient is -5 The last term, "the constant", is +4 Step-1 : Multiply the coefficient of the first term by the constant 1 • 4 = 4 Step-2 : Find two factors of 4 whose sum equals the coefficient of the middle term, which is -5 -4 + -1 = -5 That's it Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, (-4) and (-1) x4 - 4x2 - 1x2 - 4 Step-4 : Add up the first 2 terms, pulling out like factors : x2 • (x2-4) Add up the last 2 terms, pulling out common factors : 1 • (x2-4) Now add up the four terms of step 3 : (x2-1) • (x2-4)

OpenStudy (anonymous):

But we are far from done, take the first half (x2-1) Factoring: x2-1 Theory : A difference of two perfect squares, A2 - B2 can be factored into (A+B) • (A-B) Proof : (A+B) • (A-B) = A2 - AB + BA - B2 = A2 - AB + AB - B2 = A2 - B2 Note : AB = BA is the commutative property of multiplication. Note : - AB + AB equals zero and is therefore eliminated from the expression. Check : 1 is the square of 1 Check : x2 is the square of x1 Factorization is : (x + 1) • (x - 1) For the second half (x^2-4) Factoring: x2 - 4 Check : 4 is the square of 2 Check : x2 is the square of x1 Factorization is : (x + 2) • (x - 2) Now add them together to get (x + 1) • (x - 1) • (x + 2) • (x - 2)

OpenStudy (anonymous):

Now solve each single variable equation to get 4 solutions. Post if you want to see if they are correct :)

OpenStudy (anonymous):

@Nexxus ??? Do want me to double-check your work and any q's?